# RobotCycle: Assessing Cycling Safety in Urban Environments

### v1

@@ V Data Collection
VI Data Analysis
To the best of our knowledge, this is the first project capturing data from the cyclists’ perspective using a multi-sensing backpack. Our sensing platform described in section IV provides a 360-degree FOV, offering a comprehensive perspective of the cyclist’s surroundings, capturing dynamic and static elements in the routes of diverse traffic volumes and infrastructure designed in section V. This section describes the analysis we will perform on these data to assess, predict, and monitor the cyclists’ safety on the road, as indicated in our objective LABEL:objective.benchmark. To this end, we are focusing on building HD Maps, data annotation, and traffic models, which will be used in conjunction to calculate a safety score that will holistically comprehend risks associated with traffic events, allowing us to address safety concerns effectively and develop strategies to mitigate them. In this analysis, HD Maps are central as they are a fine-grained representation of the environment, which we can augment with semantic and road quality information (smoothness, potholes, etc.) – retrieved from IMU data – and dynamic agents locations. By correlating those with traffic and trajectory models we can assess the safety of the road network. Preliminary results are shown here on a test dataset we recorded with the proposed sensing platform at the Culham Science Center in Abingdon (UK) – here culham – and the uniD dataset [21].
VI-A HD Maps
HD Maps are invaluable for providing detailed contextual information about the environment, including precise lane locations, road boundaries, curbs, and landmarks. In addition, they capture local and global morphology, such as surface and geometry, detailing static infrastructure. This level of accuracy offers insights into complex intersections and infrastructure, aiding in localising all road agents and allowing us to correlate their behaviours to traffic flows and interactions. A preliminary example of an HD Map applied to culham is depicted in fig. 5.
@@ VI-B Data Annotation
To analyse the data and extract trajectory and behavioural patterns, we need to detect and track semantics of interest in the dataset. We will annotate all dynamic agents and road-layout-specific classes in the dataset, extending existing class definitions – e.g. the extensively-used CityScapes’ [22] – including infrastructural elements critical to a thorough description of the scene. Table II summarises our semantic class definition of interest, and fig. 6 shows labelled images in our culham dataset.
@@
ground
road, pavement, cycle-lane, bus-lane, parking-spot, bus-stop, lane-marking, roundabout, crossing, traffic-island, other
structure
building, fence, wall, barrier, stop-shelter, road-works, overhead-bridge, other
nature
vegetation, trunk
object
pushchair, sign, fence, pole, traffic-cone, traffic-light, wheelchair, other
vehicle
car, truck, bicycle, motorcycle, e-scooter, mobility-scooter, van, bus, on-rails, other
animate-agent
animal, pedestrian, rider, other
TABLE II: Target categories and derived subcategories for semantic labelling of image and pointcloud data focusing on infrastructural elements and dynamic agents.
@@ VI-C Traffic and Trajectory Models
Since this dataset will be taken from the perspective of a VRU instead of a vehicle, it will contain many VRU-to-VRU and VRU-to-vehicle interactions. These interactions have a higher degree of variability than interactions between vehicles, as VRUs do not necessarily follow allocated lanes, traffic lights, and road rules. Cyclists especially operate at high speeds, posing a danger to other types of road users, making predicting cyclist trajectories crucial for safety assessment.
We have begun developing these processes by leveraging the uniD dataset [21]. Figure 7 shows results for road agents’ tracked directions and velocity, divided by kind. As aforementioned, it is visible how VRUs tend to occupy more of the road surface and with more varied behaviours.
@@ VII Future Work and Open Issues
As the outlined work packages have provided encouraging preliminary results, further refinement and analysis is imperative for the project’s success. We aim to conduct a data-driven safety analysis leveraging AV research’s behavioural modelling and trajectory prediction methods. In addition, an in-depth causality analysis is required to identify each captured event’s intricate causes and effects. Our data collection is limited to private and constrained environments, capturing very few driving interactions; in the final iteration of our project, we will collect the Oxford dataset described in this paper, which we commit to making publicly available. Using this novel dataset, we aim to develop a generalisable approach to assessing safety, which, based on infrastructure, traffic volumes, and historical data, can generate insights into potentially unsafe situations, even on roads not traversed within the dataset.
@@ VIII Conclusion
This paper presents our RobotCycle project, aiming to leverage AV research and technologies to assess cycling infrastructure safety through data-driven approaches. Here, we have described the development process to build a wearable data collection device, the design of the data collection process, and the data analysis rationale. Our primary focus has been on building and testing the sensing backpack to ensure that it satisfies all design requirements so that we can achieve the project’s objectives. Initial results demonstrate the device’s ability to capture diverse and rich datasets, which are then used to extract semantic elements, HD Map, and traffic patterns. The RobotCycle project is an innovative undertaking with a mission to enhance the safety and sustainability of urban cities. We aim to generate valuable insights into the ongoing discourse on urban development, setting a benchmark for further advancements in the field.

### v3

@@ VI-B Data Annotation
To analyse the data and extract trajectory and behavioural patterns, we need to detect and track semantics of interest in the dataset. We will annotate all dynamic agents and road-layout-specific classes in the dataset, extending existing class definitions – e.g. the extensively-used CityScapes’ [25] – including infrastructure elements critical to a thorough description of the scene. Table II summarises our semantic class definition of interest and Figure 7 depicts labelled images in our culham dataset.
@@
ground road, pavement, cycle-lane, bus-lane, parking-spot, bus-stop, lane-marking, roundabout, crossing, traffic-island, other
structure building, fence, wall, barrier, stop-shelter, road-works, overhead-bridge, other
nature vegetation, trunk
object pushchair, sign, fence, pole, traffic-cone, traffic-light, wheelchair, other
vehicle car, truck, bicycle, motorcycle, e-scooter, mobility-scooter, van, bus, on-rails, other
animate-agent animal, pedestrian, rider, other
TABLE II: Target categories and derived subcategories for semantic labelling of image and pointcloud data focusing on infrastructure elements and dynamic agents.
@@ VI-C Traffic and Trajectory Models
Since this dataset will be taken from the perspective of a VRU instead of a vehicle, it will contain diverse VRU-to-VRU and VRU-to-Vehicle interactions. These interactions have higher variability as VRUs do not necessarily follow allocated lanes, traffic lights, and road rules. Especially cyclists, because they operate at higher speeds, pose a danger to other types of road users, making predicting cyclist trajectories crucial for a comprehensive safety assessment.
We have begun developing these processes by leveraging the uniD dataset [24]. Figure 8 shows results for road agents’ tracked directions and velocity, divided by kind. As expected, VRUs tend to occupy a larger portion of the road surface exhibiting a wider range of behaviours.
@@ VII Future Work and Open Issues
As the outlined work packages have provided encouraging preliminary results, further refinement and analysis are imperative for the project’s success. We aim to conduct a data-driven safety analysis leveraging behavioural modelling and trajectory prediction methods from AV research. In addition, an in-depth causality analysis is required to identify the intricate causes and effects of each event captured.
Our initial data collection is limited to private and constrained environments, capturing very few driving interactions; in the final iteration of our project, we will collect the Oxford dataset described in this paper, which we commit to making publicly available. Using this novel dataset, we aim to develop a generalisable approach to assessing safety, which, based on infrastructure, traffic volumes, and historical data, can generate insights into potentially unsafe situations, even on roads not traversed within the dataset. Within the scope of this project, our focus lies on urban environments. However, an extension to off-road areas can provide valuable insights into the behaviour of VRUs across more diverse and challenging environments [26].
@@ VIII Conclusion
This paper introduces the RobotCycle project which aims to evaluate the safety of cycling infrastructure by leveraging advancements in AV research and technologies. Here, we have described the development process of a wearable data collection device, the design of the data collection process, and the data analysis rationale. Our primary focus has been on building and testing the sensing backpack to ensure that it satisfies all design requirements so that we can achieve the project’s objectives. Initial results demonstrate the device’s ability to capture diverse and rich datasets, which are then used to extract semantic elements, HD Maps, and traffic patterns. The RobotCycle project is an innovative undertaking with a mission to enhance the safety and sustainability of urban cities. We aim to generate valuable insights into the ongoing discourse on urban development, setting a benchmark for further advancements in the field.